# Memorization to Generalization: Emergence of Diffusion Models from Associative Memory

### v1

@@ Abstract
Hopfield networks are associative memory (AM) systems, designed for storing and retrieving patterns as local minima of an energy landscape. In the classical Hopfield model, an interesting phenomenon occurs when the amount of training data reaches its critical memory load — spurious states, or unintended stable points, emerge at the end of the retrieval dynamics, leading to incorrect recall. In this work, we examine diffusion models, commonly used in generative modeling, from the perspective of AMs. The training phase of diffusion model is conceptualized as memory encoding (training data is stored in the memory). The generation phase is viewed as an attempt of memory retrieval. In the small data regime the diffusion model exhibits a strong memorization phase, where the network creates distinct basins of attraction around each sample in the training set, akin to the Hopfield model below the critical memory load. In the large data regime, a different phase appears where an increase in the size of the training set fosters the creation of new attractor states that correspond to manifolds of the generated samples. Spurious states appear at the boundary of this transition and correspond to emergent attractor states, which are absent in the training set, but, at the same time, have distinct basins of attraction around them. Our findings provide: a novel perspective on the memorization-generalization phenomenon in diffusion models via the lens of AMs, theoretical prediction of existence of spurious states, empirical validation of this prediction in commonly-used diffusion models. denotes equal contribution.
@@ Appendix D Experimental Details on the Transition Mapping
Figure 9: The fractions of memorized, spurious, and generalized samples in synthetic sets across different training sizes for CIFAR10 and various model’s widths. As the training data size $K$ increases, memorization decreases while the emergence of spurious patterns follows and the fraction of generalized patterns steadily increase (see top row). The fraction of spurious patterns rises and decreases at the boundary between the memorization and generalization phases (see bottom row). Additionally, with a smaller number of parameters, the memorization-generalization transition happens at an earlier stage (or smaller $K$), since the model has a memorization capacity. In contrast, as the model’s size increases, the transition happens at an later stage (or larger $K$), where it is able to memorize more training data. The numbering, e.g., $64$, $96$, and $128$, denotes the initial latent dimension of the U-Net, see Table 3.
@@
Additional Experiments. Inspired by [45, 46], we extend our investigation on the memorization-generalization transition by studying the effects of varying the parameter size of the diffusion model for CIFAR10. The results of this experiment are recorded in Figure 9. Specifically, we trained two additional sets of 38 diffusion models, with the initial latent dimensions of 64 and 96 respectively, and repeat the same experimental setup delineated above, see also Table 3. For U-Net (64), we identified the critical points, $A$ and $B$, to be $500$ and $8000$ respectively. Meanwhile, for U-Net (96), these critical points are $2000$ and $16000$. To select $\delta_{m}$ and $\delta_{s}$, we followed the same process detailed above by observing the distance histograms to see where bimodality occurs with respect to the distance value, and manually inspecting the images to ensure there is a minimal amount of false positives in memorized and spurious samples. For U-Net (64), the threshold values $\delta_{m}$ and $\delta_{s}$ are both selected to be $0.02$ — while for U-Net (96), the threshold values $\delta_{m}$ and $\delta_{s}$ are chosen as $0.03$ and $0.02$, respectively.
@@
With the smaller U-Net (64) model, we observe the memorization-generalization transition happens at an earlier stage with respect to the larger U-Net (128) model, where the peak of spurious samples appears at $K=2568$ instead of $K=5862$. Similarly, for U-Net (96), the peak of spurious samples happens at an early stage of $K=3931$ instead. Overall, these results indicate that as the parameter size of the diffusion model increases, the memorization capacity of the model is increased and thus, generalization is delayed since the model can memorize more of its training data denoted by the appearance of spurious patterns happening at a larger training data size $K$. In contrast, when the memorization capacity is reduced, generalization is initiated at an earlier stage (or smaller training data size) with respect to when the parameter size of the model is larger. Most importantly, the emergence of spurious patterns remains consistent despite the variations in the parameter size of the diffusion model.
@@
Hardware Details. Training of the models detailed in Table 3 were done using many NVIDIA Tesla V100 GPUs (a single GPU was used to train each model). Each GPU has 32GB of memory and is linked with IBM Power9 processors which clocked at 3.15 GHz. Each CIFAR10 model took roughly 18, 20, and 22 hours to train for different initial latent dimensions, e.g., 64, 96, 128, respectively. Meanwhile, each LSUN-CHURCH model requires roughly 56 hours of training. In contrast, FASHION-MNIST and MNIST models require approximately 17 hours to be fully trained (per model). Please note that these durations are with respect to the required training iterations for each model and their corresponding dataset detailed in Table 3. Please note that we utilized the same hardware for our experiments as well.

### v2

@@ Abstract
Hopfield networks are associative memory (AM) systems, designed for storing and retrieving patterns as local minima of an energy landscape. In the classical Hopfield model, an interesting phenomenon occurs when the amount of training data reaches its critical memory load — spurious states, or unintended stable points, emerge at the end of the retrieval dynamics, leading to incorrect recall. In this work, we examine diffusion models, commonly used in generative modeling, from the perspective of AMs. The training phase of diffusion model is conceptualized as memory encoding (training data is stored in the memory). The generation phase is viewed as an attempt of memory retrieval. In the small data regime the diffusion model exhibits a strong memorization phase, where the network creates distinct basins of attraction around each sample in the training set, akin to the Hopfield model below the critical memory load. In the large data regime, a different phase appears where an increase in the size of the training set fosters the creation of new attractor states that correspond to manifolds of the generated samples. Spurious states appear at the boundary of this transition and correspond to emergent attractor states, which are absent in the training set, but, at the same time, have distinct basins of attraction around them. Our findings provide: a novel perspective on the memorization-generalization phenomenon in diffusion models via the lens of AMs, theoretical prediction of existence of spurious states, empirical validation of this prediction in commonly-used diffusion models.The code is available at: https://github.com/Lemon-cmd/Diffusion-Models-and-Associative-Memory denotes equal contribution.
@@ Appendix D Experimental Details on the Transition Mapping
Figure 9: The fractions of memorized, spurious, and generalized samples in synthetic sets across different training sizes for CIFAR10 [63] and various model’s widths. As the training data size $K$ increases, memorization decreases while the emergence of spurious patterns follows and the fraction of generalized patterns steadily increase (see top row). The fraction of spurious patterns rises and decreases at the boundary between the memorization and generalization phases (see bottom row). Additionally, with a smaller number of parameters, the memorization-generalization transition happens at an earlier stage (or smaller $K$), since the model has a memorization capacity. In contrast, as the model’s size increases, the transition happens at an later stage (or larger $K$), where it is able to memorize more training data. The numbering, e.g., $64$, $96$, and $128$, denotes the initial latent dimension of the U-Net, see Table 3.
@@
Additional Experiments. Inspired by [45, 46], we extend our investigation on the memorization-generalization transition by studying the effects of varying the parameter size of the diffusion model for CIFAR10 [63] and LSUN-CHURCH [66]. The results of this experiment are recorded in Figs. (9) and (10), respectively. Specifically, we trained two additional sets of 38 diffusion models, with the initial latent dimensions of 64 and 96 respectively, and repeat the same experimental setup delineated above, see also Table 3.
For CIFAR10 and the U-Net (64) models, we identified the critical points, $A$ and $B$, to be $500$ and $8000$ respectively. Meanwhile, for the U-Net (96) models, these critical points are $2000$ and $16000$. To select $\delta_{m}$ and $\delta_{s}$, we followed the same process detailed above by observing the distance histograms to see where bimodality occurs with respect to the distance value, and manually inspecting the images to ensure there is a minimal amount of false positives in memorized and spurious samples. For U-Net (64) models, the threshold values $\delta_{m}$ and $\delta_{s}$ are both selected to be $0.02$ based on the bimodality of the distance histograms observed at $K=2568$ — meanwhile, for U-Net (96), the threshold values $\delta_{m}$ and $\delta_{s}$ are chosen as $0.03$ and $0.02$ according to distance histograms observed at $K=3931$.
Similarly, for the LSUN-CHURCH models, we identified the critical points, $A$ and $B$, to be $1000$ and $8000$ for the U-Net (64) models; and for the larger U-Net (128) models, $4000$ and $16000$ as the critical points. The threshold values, $\delta_{m}$ and $\delta_{s}$, are both selected to be $0.04$ after observing the bimodality of the distances at $K=3413$ — meanwhile, for U-Net (128), these values are selected to be $\delta_{s}=0.04$ and $\delta_{m}=0.1$ after observing the distance histograms at $K=5241$.
Figure 10: The fractions of memorized, spurious, and generalized samples in synthetic sets across different training sizes for LSUN-CHURCH [66] and various model’s widths. As the training data size $K$ increases, memorization decreases while the emergence of spurious patterns follows and the fraction of generalized patterns steadily increase (see top row). The fraction of spurious patterns rises and decreases at the boundary between the memorization and generalization phases (see bottom row). Additionally, with a smaller number of parameters, the memorization-generalization transition happens at an earlier stage (or smaller $K$), since the model has a memorization capacity. In contrast, as the model’s size increases, the transition happens at an later stage (or larger $K$), where it is able to memorize more training data. The numbering, e.g., $64$, $96$, and $128$, denotes the initial latent dimension of the U-Net, see Table 3.
@@
With the smaller U-Net (64) models, we observe the memorization-generalization transition happens at an earlier stage with respect to the larger U-Net (128) models, where the peak of spurious samples appears at $K=2568$ instead of $K=5862$ for CIFAR10 and $K=1000$ instead of $K=4000$ for LSUN-CHURCH. Similarly, for CIFAR10 and its U-Net (96) models, the peak of spurious samples happens at an early stage of $K=3931$ instead. Overall, these results indicate that as the parameter size of the diffusion model increases, the memorization capacity of the model is increased and thus, generalization is delayed since the model can memorize more of its training data denoted by the appearance of spurious patterns happening at a larger training data size $K$. In contrast, when the memorization capacity is reduced, generalization is initiated at an earlier stage (or smaller training data size) with respect to when the parameter size of the model is larger. Most importantly, the emergence of spurious patterns remains consistent despite the variations in the parameter size of the diffusion model.
@@
Hardware Details. The training of the models detailed in Table 3 was done using NVIDIA Tesla V100 GPUs, where a single GPU was used to train each model. Each GPU has 32GB of memory and is linked with IBM Power9 processors which clocked at 3.15 GHz. Each CIFAR10 model took roughly 18, 20, and 22 hours to train for different initial latent dimensions, e.g., 64, 96, 128, respectively. Meanwhile, each LSUN-CHURCH model requires roughly 48, 56, and 103 hours of training with respect to the model’s initial latent dimension. In contrast, FASHION-MNIST and MNIST models require approximately 17 hours to be fully trained (per model). Please note that these durations are with respect to the required training iterations for each model and their corresponding dataset detailed in Table 3. Please note that we utilized the same hardware for our experiments as well.